OpenStudy (anonymous):

Delip measured the speeds, x km per hour, of 70 cars on a road where the speed limit is 60 km per hour. His results are summarised by Σ(x − 60) = 245. His friend Sachim used values of (x − 50) to calculate the mean. (ii) Find Σ(x−50).

OpenStudy (anonymous):

@phi

OpenStudy (experimentx):

mean = 245/70+60 mean = y/70 + 50 y = 245+700 ... though no so sure

OpenStudy (phi):

70 cars, Σ(x − 60) = 245 Σx - 60*70=245 Σx = 245+60*70 Σx/70 = x_mean = 60+3.5= 63.5 (x − 50) to calculate the mean. I assume he found the mean of x-50 (Σ(x-50))/70 = Σx/70 - 50*70/70 = x_mean-50 (Σ(x-50))/70 = 63.5-50 = 13.5 Σ(x-50)= 70*13.5 Σ(x-50)= 945

OpenStudy (phi):

It does match experimentX's result

OpenStudy (experimentx):

usually my hunch is right ... thanks for confirming